OpenStudy (anonymous):

if ∀x were false for all x , will it be true ? for example: ∀x Q(x) where Q(x) = ,,, x=x^2 and the domain consist of integers beside 0 and 1.

OpenStudy (anonymous):

also here's a question: ∃x(n=-n) and the domain consist of integers , find the truth value well it's false for every integer but what about 0 ? i know a 0 can't be negative but it will be the same as (n=n) ?

OpenStudy (theeric):

Hi! Math proofs is really cool, but it seems to take a special mode of thinking! An axiomatic way of thinking. I'm just okay at proving. But I might be able to help. In your first question, I don't understand your notation. Sorry. = ,,, x= what does that mean? With your second question, \(n=-n=(-1)n\) So, consider \(n=0\). Let me prove this... \(n=0\) By multiplying by \(-1\), we see that \((-1)n=(-1)(0)\) \(\implies-n=0\) Now \(n=-n\) where \(n=0\) Or you can just say \(n=0=(-1)0=(-1)n=-n\) \(\implies n=-n\)

OpenStudy (theeric):

For that second question, both ways are just as valid. I think.

OpenStudy (anonymous):

I mean that Q(x) is (x=x^2)

OpenStudy (theeric):

Oh! There's a name for that \(x=x^2\) relationship! I forget what it is... I think it starts with an A, but I could be wrong.

OpenStudy (anonymous):

ah thank you man

OpenStudy (theeric):

So, here's the first question (restated) For all \(x\) where \(x\) can be any integer except for \(0\) and \(1\), does \(x=x^2\)? If it does, then \(x=x^2\) is true for that domain. If not, \(x=x^2\) is false. And I do mean ALL \(x\) must be true when we say "all \(x\)." Or \(\forall x\), if you prefer.

OpenStudy (theeric):

For the second problem there, you also have to determine truth or falsity, but for each integer separately. I think you have the idea. Zero is the only number that makes it true,j and 0 is an integer, so it's in the domain. I can prove that all other numbers won't work, but this problem doesn't ask for that.

OpenStudy (theeric):

Let me know if you have any questions.

OpenStudy (anonymous):

I think about the first question if it was all false then it's true , since all of x is the same

OpenStudy (anonymous):

@theEric are you there ? :D what do you think ? I think if " my first question " is false for all the domain and you see that 0 and 1 are not included so it's false for all x , and my proof said that if it's true for all x then it's true since we're talking about all x and if it's false for all x then it's also true.

OpenStudy (theeric):

The way I see it, is that it's a typical "TRUE or FALSE" statement type of question. Like the kind I usually saw in history class. Anyway. Now the statement is For all \(x\) in the domain (like \(\infty ,...,-2, -1, 2, 3,..., \infty\)) we see that \(x=x^2\). Do you see that for all \(x\) in the domain?

OpenStudy (anonymous):

it's false for all x

OpenStudy (theeric):

It is false for all \(x\)!

OpenStudy (theeric):

\(x=x^2\) is false for all \(x\)!

OpenStudy (theeric):

So, the statement is false in it's entirety.

OpenStudy (anonymous):

and this is my question , if we're asked to give the truth value of ∀x for this one , will it be true or false ?

OpenStudy (theeric):

I say false. It can be true ONLY IF ALL \(x\) in the domain make it true. Turns out no \(x\) makes it true. So, the statement is false.

OpenStudy (anonymous):

thank you ^^

OpenStudy (theeric):

You're welcome!

OpenStudy (theeric):

Good luck with any other problems that you have to do!